OpenStudy (anonymous):

evaluate 7^35*7^-21*7^-8

OpenStudy (camerondoherty):

117,649 i think

OpenStudy (anonymous):

I agree

OpenStudy (shamil98):

\[a^b*a^c = a^{b+c}\] \[7^{35}*7^{-21}*7^{-8} = 7^{35 + (-21) + (-8)} = 7^6\]

OpenStudy (shamil98):

When evaluated it is 117,649, but the process is good to know.

OpenStudy (anonymous):

thank you guys, I was confused but I understand now. Free fan for all of you :)

OpenStudy (camerondoherty):

lol

OpenStudy (camerondoherty):

medals r kool too tho ;)